# Foundation Models for Atomistic Simulation of Chemistry and Materials

### v1

@@ Abstract
corresponding authors: aditik1@berkeley.edu, bingqingcheng@berkeley.edu, smblau@lbl.gov, thg@berkeley.edu
Glossary
  • Atomic Cluster Expansion: A framework for constructing atom-centered descriptors using systematically improvable body-ordered terms, which helps in describing atomic environments in machine learning interatomic potentials.
  • Attention: A mechanism used in state-of-the-art machine learning models that dynamically weighs different input elements based on their relevance to a given task, enabling models to focus on the most important information.
  • Born-Oppenheimer Approximation: An approximation in quantum chemistry that separates the motion of nuclei and electrons in a molecule, simplifying the Schrödinger equation.
  • Compute Scaling: The ability of a system to handle more users, data, or load without losing performance.
  • Coupled Cluster: A method in computational chemistry to describe the electronic structure of many-electron systems, often regarded as the gold-standard of accuracy in quantum chemistry.
  • Density Functional Theory (DFT): A method in computational chemistry to describe the electronic structure of many-electron systems, often regarded as an affordable and sufficiently accurate choice in quantum chemistry.
  • Direct Force Model: A model that directly predicts atomic force vectors without computing the gradient of a scalar potential energy, thereby accelerating model inference while sacrificing their connection with the energy and the zero-curl property necessary for energy-conserving simulations.
  • Equivariance: A property of a model where its predictions and/or representations undergo the same changes under transformations, in particular rotations, of the input data, important for ensuring physical consistency in simulations.
  • Fine-Tuning: A process where a pre-trained model is further trained, often on a labeled dataset with task-specific supervision, to improve its performance on a target application.
  • Foundation Model (FM): A large-scale model pre-trained on diverse data to capture complex patterns, which can be fine-tuned for various downstream tasks.
  • Gradient-Based Force Model: A model that predicts atomic force vectors as the gradient of a scalar potential energy, thereby guaranteeing the conservation of energy in chemical simulations.
  • Graph Neural Network (GNN): A type of neural network designed to process data structured as graphs, where nodes represent entities and edges represent relationships between them.
  • Scaling Laws: Empirical rules that describe how the performance of a model improves with increasing model size, training data, and computational resources.
  • Inference: The use of a machine learning model to make predictions on new data not seen during training.
  • Invariance: A property of a model where its predictions and/or representations remain unchanged under transformations, in particular translations and rotations, of the input data, important for ensuring physical consistency in simulations.
  • Machine Learning Interatomic Potential (MLIP): A model that uses machine learning techniques to predict the potential energy surface of atomic systems, enabling simulations of molecular dynamics, geometry optimizations, and other downstream applications.
  • Message Passing: A process in graph neural networks where information is exchanged between nodes to capture relationships and interactions in the data.
  • Model Distillation: A technique where a simpler model is trained to replicate the behavior of a larger, more complex model, transferring knowledge from the teacher to the student model.
  • Potential Energy Surface: A property in computational chemistry that defines the electronic energy as a function of nuclear positions under the Born-Oppenheimer approximation.
  • Pre-Training: A stage in machine learning where a model is trained on a large dataset to learn general-purpose representations.
  • Self-Supervised Learning: A type of machine learning where models are trained on unlabeled data by predicting parts of the input from other parts, often used for pre-training large models.
  • Supervised Learning: A type of machine learning where models are trained on labeled data by predicting a pre-calculated output from the input.
  • Training: The process by which the parameters of a machine learning model are learned, typically via gradient-based optimization of an objective function averaged over a large quantity of data.
  • Transfer Learning: A machine learning technique where a model developed for one task is reused as the starting point for a model on another distinct but relevant task.
  • Uncertainty Quantification: Techniques used in machine learning to estimate confidence in model predictions, helping to assess reliability and robustness in decision-making.
  • Universal Potential: A machine learning model trained to predict potential energies and forces across a wide range of chemical systems instead of a single specific system.
@@ 1 Introduction
The fundamental principle underlying the description of atomistic chemistry and materials science is the Schrödinger equation, whereby a chemical system and its physical properties can be uniquely defined by Cartesian and spin coordinates of the component atoms and system net charge. But as famously stated by Dirac after the discovery of quantum mechanics "..the underlying physical laws necessary for the whole of chemistry are thus completely known, and the difficulty lies only in the fact that the exact application of these laws leads to equations much too complicated to be soluble." This has led to a series of well-controlled approximations, such as the Born-Oppenheimer separation of nuclei and electronic degrees of freedom, and the transformation of the Schrödinger eigenvalue equation into an algebraic framework, that has allowed for good model chemistries to be solved for systems beyond the hydrogen molecule.[1] Even so, the most widely used approximate electronic structure methods based on Density Functional Theory (DFT)[2] remain largely inaccessible for long timescales and large length scales and system sizes that define many of the interesting areas of chemistry and materials science.
@@
While LLMs and LVMs themselves hold great potential to impact chemistry and materials science[5, 6, 7, 8], in this perspective we are focused on whether FMs can directly capture the potential energy surface (PES) via machine learned interatomic potentials (MLIPs).[9, 10] The end goal of such a model is to enable large-scale atomistic simulation and analysis, including the ability to derive any observable of interest, for any molecular or periodic system, either with a small amount of system-specific fine-tuning[11] or without any fine-tuning at all.
To achieve such a goal, this perspective first provides more concrete definitions for what defines a FM, and how a FM for atomistic simulation should be distinguished from a “universal potential” or even a more restricted definition such as transfer learning. In particular, our criteria begins with the ability of a MLIP FM to exhibit scaling laws and pre-training and data strategies that have made LLMs and LVMs so powerful. With this context, we examine the capacity of current model architectures and identify limitations and opportunities that are relevant to their scalability to larger learnable parameters in order to become more expressive and foundational to increasing amounts of data. We also consider what type and amount of data of scale will be required for a generally applicable pre-trained FM, how that compares to the data that already exists, and where we expect fine-tuning data to be most important for chemistry and materials. We further examine the importance of more advanced training strategies, including self-supervised and unsupervised pre-training as well as model distillation, and highlight that new methods for physical infusion need to be relevant and feasible when training at scale on advanced hardware. Finally, we consider how the performance of MLIPs is currently evaluated, and to suggest new benchmarks for demonstrating scaling laws and the ability to be broadly fine-tuned for a wide range of downstream applications, which is particularly important given the breadth of desired observables in chemistry and materials.
@@ 2 Characteristics of Foundation Models
One important clarification that we hope to address in this perspective is to distinguish the term “foundation model" from other related terms such as “transferable” or “universal”, particularly in the context of MLIPs. The popular MLIP MACE-MP-0[24] is trained on data from the Materials Project[25, 26] generated at the PBE level of DFT. The resulting ML model exhibits good accuracy in energy and force predictions against PBE on other crystalline unit cells and exhibits excellent stability in the numerical solution to Hamilton’s equation of motion across related chemical and materials systems. This has opened up a whole new range of applications, including free energy calculations, simulating spectroscopic observables, and explorations of phase diagrams. However, we would stipulate that MACE-MP-0 is not a foundation model, but an excellent universal PBE potential, given that it is trained with supervision to do a very specific task: predict energy and force labels for a single chemical domain at one level of theory. Further, it has not yet demonstrated fine-tuning performance across a broad range of down-stream tasks superior to task-specific models trained from-scratch. Another key limitation is that MACE-MP-0[24], like many universal MLIPs, does not attempt to distinguish between systems with different total charge and spin, whose energy and forces would be different although the underlying input structure is the same. At present MLIP FMs do not show compliance with heuristic scaling laws nor have demonstrated that a large-scale model is ready for emergent predictions, such as predicting higher quality CCSD(T) data or field-dependent properties, for example.
@@ 3.1 Machine Learning Architectures for Chemistry and Materials
The early Behler-Parrinello neural network potentials[27] used hand-crafted two-body and three-body descriptors for describing atomic environments. Later, DeePMD models[28] automated the optimization and discovery of such descriptors, reducing reliance on manual feature engineering. The atomic cluster expansion (ACE) method[29] introduced a unified and generalizable framework for constructing atom-centered descriptors using systematically improvable body-ordered terms. Graph neural networks (GNNs) then broadened the landscape, allowing for flexible and expressive architectures that learn features by iteratively exchanging information between neighboring atoms, thus capturing many-body effects[30, 31, 32]. More recently, GNN models have incorporated different forms of graph attention mechanisms. [33, 34]. These models can be viewed as a more general way of performing message passing between atoms.
@@
Figure 1: (a) Distribution of softening scales ($f_{\text{MLIP}}/f_{\text{DFT}}$) sampled from 1,000 high-energy configurations using the structures from the WBM dataset.[47] (Reproduced from Ref. [48]). (b) An example of two charged molecules that are far apart and do not interact due to (c) short-range implicit bias from GNNs and message passing, and (d) a new ML model that incorporates long-ranged information [49]. (e) Depiction of instability in molecular simulations with MLIPs, which can occur even at very small time discretizations [50].
There are possible issues of the current MLIP architectures that should be examined more closely. The first is transferability outside the pre-training data. While recently developed universal MLIPs demonstrate reasonable performance in predicting materials stability on element-wise out-of-distribution datasets[47, 51], these evaluations primarily focus on local energy minimum searches through structure relaxation of crystalline materials, sharing similar prototypes in the Materials Project database. A systematic softening (underestimation) of the ML-predicted PESs and interatomic forces[52] compared to DFT ground truth was observed in a recent benchmark study as seen in Figure 1a. [48] The underestimation of the PES is particularly significant for high-energy atomic configurations, which limits its application to molecular dynamics simulations and rare-event samplings. Recent large data efforts[53, 54] and open-sourced industry-level universal MLIPs[55] that include more off-equilibrium atomic configurations show promise in addressing these challenges as discussed in Section 3.2. However, direct transferability of universal MLIPs for specific downstream molecular or materials modeling remains uncertain, as they are not yet demonstrably better than models built directly for the new system or property from scratch. We return to this issue in Section 3.4.
Another limitation is the lack of incorporation of long-range interactions. Although short-range potentials may be sufficient to describe most properties of homogeneous bulk systems [56], they may fail for interfaces [57], dielectric response [58, 59], and dilute ionic solutions with Debye-Hückel screening. For example, consider two charged molecules that are separated by a distance greater than the cutoff radius used in most MLIPs (Fig.1b). These models, which focus on short-range interactions, will fail to accurately describe the electrostatic forces between the molecules (Fig.1c). This issue cannot be resolved simply by using message passing, as the molecules exist on separate graphs and do not interact within the framework of short-range biases inherent to message-passing neural networks. On the other hand, MLIPs with built-in long-range corrections [60, 49] can effectively solve such artifacts (Fig.1d).
@@ 3.2 Data Requirements, Pitfalls, and Challenges
ANI-1x[72]/ANI-2x[68] 8.9M $\omega$B97X/6-31G* Dimer scan
@@
Alexandria[80] 30M PBE(+U)/PAW Relaxation trajectory
@@
Regarding elemental diversity, most plane-wave DFT datasets follow the standards set by the Materials Project[25] and thus cover most of the periodic table up through plutonium. In contrast, molecular force field and DFT datasets are considerably more constrained in their elemental coverage, with most datasets only including C, H, N, and O and some coverage of S, F, Cl, and Br. The only molecular datasets of over 1 million data points with both energy and forces labels and which go beyond this set of elements are SPICE[76] (P, Li, Na, Mg, K, Ca, and I) and AIMNet2 (P, I, Si, B, As, and Se). The only large-scale datasets with variable-charge species are SPICE(2)[76], AIMNet2[67], AIMNet-NSE[83], and solvated protein fragments[74]/GEMS[75], while AIMNet-NSE[83] is a dataset with variable spin species (although it carries no force labels). Another consideration for chemical diversity is the variety of chemical interactions captured by the data. While most of the molecular datasets mentioned thus far focus on isolated molecules, SPICE[76] makes an effort to capture varied intermolecular interactions between molecules, including some metal ions, solvated protein fragments[74]/GEMS[75] includes protein fragments with multiple nearby solvent species, and ANI-2x[68] includes dimer scans and water clusters.
@@
In order to create structural diversity, large-scale data-acquisition efforts use a collection of sampling techniques such as normal modes, ab initio and empirical molecular dynamics, minimization and relaxation, reaction path sampling, metadynamics, and active learning, as outlined in Table 1. Structural diversity also encompasses the nature of the regions of the PES covered by the dataset. Having a broad range of forces between 0 eV/Å and $\sim$10 eV/Å is important to effectively learn chemically-critical components of the PES encountered during MD, geometry optimization, or reaction path and transition state optimization. To that end, Transition-1x is the only large-scale dataset which explicitly samples reaction paths and transition states – critical areas for any truly foundational MLIP.[84, 85] Extremely high forces (>50 eV/Å) can be detrimental to training and are usually filtered out, although high-energy and force data can be weighted with a Boltzmann factor[86], which still provides the model important chemical information about these regions. Finally, given that LLM FMs are pre-trained in an unsupervised manner on unlabeled data, it is worth asking what sort of data would be required to realize an analogous procedure in the chemical context. Ji and co-workers recently demonstrated that self-supervised pre-training on 800 million unlabeled 3D conformations of organic molecules improved down-stream performance on QM9 property prediction.[64]
The specific areas where we believe the community should devote major effort to large-scale data generation are shown in Figure 2. For molecular DFT, the majority of the periodic table remains poorly covered, and under-representation of metal-organic complexes, complex electrolytes, metalloenzymes, and disordered proteins underscores the need for substantial dataset generation efforts beyond organic elements (Fig. 2a). These types of systems could also address three other major gaps: non-covalent interactions (Fig. 2b), systems where long-range interactions are critical (Fig. 2e), and variable charge and spin states (Fig. 2c) which have been poorly covered by current datasets. We suggest that, beyond open-shell metals, combustion chemistry and atmospheric chemistry are spaces where different spin states are important and could be valuable. The final major gap in the molecular space is reactivity, particularly involving complex mechanisms and reactions in the condensed phase (Fig. 2d).
In the context of plane-wave DFT, we believe that major gaps include solid-solid interfaces (Fig. 2f), solid-liquid interfaces (Fig. 2i), surfaces under different applied potentials, (Fig. 2g), and defects (Fig. 2h). We note that some of these areas will require the use of better density functionals beyond the ubiquitous PBE and RPBE. Analogous to spin states for molecular properties, additional data generation for magnetic properties could also aid in correct magnetic moment initialization for diverse materials.
@@
Figure 2: Proposed areas where substantial additional high-quality DFT data should be generated in order to realize a general and transferable foundational MLIP. a) Beyond-organic elements. b) Non-covalent interactions. c) variable charge and spin states. d) Complex reactivity. e) long-range interactions. f) Solid-solid interfaces. g) Surfaces under applied potential. h) Defects. i) Solid-liquid interfaces. j) Rich quantum information.
@@
Finally, we believe there is a major need for large-scale datasets of rich quantum information, such as electron densities, orbital information, QTAIM descriptors, NBO interactions, and beyond (Fig. 2j). We note that infusing rich quantum information into ML models has recently been shown to be beneficial[87, 88, 89, 90], and believe that powerful FM capabilities could be unlocked with such data. For example, MLIP FMs could be pre-trained on this extra information by swapping prediction schemes to predict atom and bond-level values. This could pave the way for foundational chemical models that go beyond MLIPs, serving as powerful representation models that can be adapted to predict forces and energies downstream.
@@
Going beyond DFT is desirable. Here it is appropriate to highlight the heroic effort of Smith and co-workers in generating the ANI-1ccx data set containing energies and forces of 500,000 organic molecules obtained with an accurate CCSD(T)*/CBS composite extrapolation.[95, 72]. This protocol has deviations from the Schrödinger limit that are larger than normally acceptable for small, very high accuracy datasets,[101] but are much smaller than the best DFT errors, even in large basis sets.[2] Continued diverse collections of gold-standard CCSD(T) data (or beyond when necessary) would be extremely valuable for fine-tuning. Another result worth highlighting is the CCSD(T) assessment[102] of over 10,000 barrier heights in RDB7, which led to unexpectedly large RMS errors from the normally reliable $\omega$B97X-D3 functional. The largest of these deviations were resolved by removing orbital instabilities and allowing the DFT to be spin-polarized when necessary.[103] However this highlights the limitations of even hybrid DFT for reactive systems where strong correlation effects are at play, as revealed by such symmetry-breaking.[104] For the most part, CC benchmarks for condensed matter periodic systems lie in the future, although there has been notable progress in CC methods for materials science.[105]
The third factor are other error sources that prevent a given “model/basis” combination from yielding a well-defined value. In particular, DFT calculations on materials require careful attention to best practices[106, 100] to ensure reproducibility, beginning with choice of cutoffs, quadrature quality, and convergence criteria. Such factors have been discussed and investigated for the OC20 data.[107] There are also non-trivial additional sources of error. Reaching the thermodynamic limit is particularly challenging for hybrid functionals.[108] Pseudopotential errors, typically not present in molecular calculations involving light elements, also require careful evaluation[109, 110, 111]. Thresholds and cutoffs also affect molecular calculations, but can be controlled, and combined with SCF algorithms that ensure descent.[112] The local correlation methods needed for CCSD(T) on large molecules[113] are also subject to domain errors.[114] They scale with size, and may impact accuracy[115, 116] and so ideally should be removed and replaced by strict numerical thresholds.[117, 118]

### v3

@@ Abstract
*corresponding authors: aditik1@berkeley.edu, bingqingcheng@berkeley.edu, smblau@lbl.gov, thg@berkeley.edu
@@ 1 Introduction
The fundamental principle underlying the description of atomistic chemistry and materials science is the Schrödinger equation, whereby a chemical system and its physical properties can be uniquely defined by Cartesian and spin coordinates of the component atoms and system net charge. But as famously stated by Dirac after the discovery of quantum mechanics "..the underlying physical laws necessary for the whole of chemistry are thus completely known, and the difficulty lies only in the fact that the exact application of these laws leads to equations much too complicated to be soluble." This has led to a series of well-controlled approximations, such as the Born-Oppenheimer separation of nuclei and electronic degrees of freedom, and the transformation of the Schrödinger eigenvalue equation into an algebraic framework, that has allowed for good model chemistries to be solved for systems beyond the hydrogen molecule.[1] Just as important is the need to satisfy the laws of statistical mechanics that requires the high dimensional integration over coordinates of Boltzmann weighted configurations to predict observables, usually collected over time trajectories with the hope that ergodicity is satisfied. Even so, the most widely used approximate electronic structure methods based on Density Functional Theory (DFT)[2] combined with ab initio molecular dynamics (AIMD) remains inaccessible for long timescales and large length scales and system sizes that define many of the interesting areas of chemistry and materials science.
@@
While LLMs and LVMs themselves hold great potential to impact chemistry and materials science[5, 6, 7, 8, 9], in this perspective we are focused on whether FMs can directly capture the potential energy surface (PES) via machine learned interatomic potentials (MLIPs).[10, 11] The end goal of such a model is to enable large-scale atomistic simulation and analysis, including the ability to derive any observable of interest, for any molecular or periodic system, either with a small amount of system-specific fine-tuning[12] or without any fine-tuning at all. Importantly, it would supersede the painful acquisition of new labeled data and the hidden expense of often hundreds of rounds of retraining of the ML model used in active learning to develop a robust MLIP.[13, 14, 15, 16]
To achieve such a goal, this perspective first provides more concrete definitions for what defines a FM, and how a FM for atomistic simulation should be distinguished from a “universal potential” or even a more restricted definition such as transfer learning. In particular, our criteria begins with the ability of a MLIP FM to exhibit scaling laws and pre-training and data strategies that have made LLMs and LVMs so powerful. With this context, we examine the capacity of current model architectures and identify limitations and opportunities that are relevant to their scalability to larger learnable parameters in order to become more expressive and foundational to increasing amounts of data. We also consider what type and amount of data of scale will be required for a generally applicable pre-trained FM, how that compares to the data that already exists, and where we expect fine-tuning data to be most important for chemistry and materials. In this perspective we focus primarily on FMs trained on energy and force data to address outstanding problems in chemistry and materials through molecular simulation. But it should be pointed out at this juncture that different modality data exists, such as images and text extraction for catalytic discovery[17], prompt engineering of LLMs for predicting synthetic pathways of small drug molecules[9], and use of video data such as that generated for liquid-phase TEM imaging[18]. We further examine the importance of more advanced training strategies, including self-supervised and unsupervised pre-training as well as model distillation, and highlight that new methods for physical infusion need to be relevant and feasible when training at scale on advanced hardware. Finally, we consider how the performance of MLIPs is currently evaluated, and to suggest new benchmarks for demonstrating scaling laws and the ability to be broadly fine-tuned for a wide range of downstream applications, which is particularly important given the breadth of desired observables in chemistry and materials.
@@ 2 Characteristics of Foundation Models
One important clarification that we hope to address in this perspective is to distinguish the term “foundation model" from other related terms such as “transferable” or “universal”, or large atomic models (LAMs)[31], particularly in the context of MLIPs. The popular MLIP MACE-MP-0[32] is trained on data from the Materials Project[33, 34] generated at the PBE level of DFT. The resulting ML model exhibits good accuracy in energy and force predictions against PBE on other crystalline unit cells and exhibits excellent stability in the numerical solution to Hamilton’s equation of motion across related chemical and materials systems. This has opened up a whole new range of applications, including free energy calculations, simulating spectroscopic observables, and explorations of phase diagrams. However, we would stipulate that MACE-MP-0 is not a foundation model, but an excellent universal PBE potential, given that it is trained with supervision to do a very specific task: predict energy and force labels for a single chemical domain at one level of theory. Another key limitation is that MACE-MP-0[32], like many universal MLIPs, does not attempt to distinguish between systems with different total charge and spin, whose energy and forces would be different although the underlying input structure is the same.
We offer that to be a true FM preferably all of the following features should be demonstrated. First is that the zero-shot or fine-tuned MLIP FM should show performance across a broad range of down-stream tasks that is superior to task-specific models trained from-scratch [31, 35]. Second, MLIP FMs would show compliance with heuristic scaling laws in regards how performance improves as model parameters and training data increases with increasing compute resources. Finally, a demonstration that a large-scale FM MLIP would have emergent capabilities, for example predicting higher quality CCSD(T) data or magnetic field-dependent properties, properties that are very different than the original underlying Cartesian geometries and DFT energy training data.
@@ 3.1 Machine Learning Architectures for Chemistry and Materials
The early Behler-Parrinello neural network potentials[36] used hand-crafted two-body and three-body descriptors for describing atomic environments. Later, DeePMD models[37] automated the optimization and discovery of such descriptors, reducing reliance on manual feature engineering. The atomic cluster expansion (ACE) method[38] introduced a unified and generalizable framework for constructing atom-centered descriptors using systematically improvable body-ordered terms. Graph neural networks (GNNs) then broadened the landscape, allowing for flexible and expressive architectures that learn features by iteratively exchanging information between neighboring atoms, thus capturing many-body effects[39, 40, 41, 42]. More recently, GNN models have incorporated different forms of graph attention mechanisms. [43, 44]. These models can be viewed as a more general way of performing message passing between atoms.
@@
There are possible issues of the current MLIP architectures that should be examined more closely. The first is transferability outside the pre-training data. While recently developed universal MLIPs demonstrate reasonable performance in predicting materials stability on element-wise out-of-distribution datasets[57, 58], these evaluations primarily focus on local energy minimum searches through structure relaxation of crystalline materials, sharing similar prototypes in the Materials Project database. A systematic softening (underestimation) of the ML-predicted PESs and interatomic forces[59] compared to DFT ground truth was observed in a recent benchmark study as seen in Figure 1a. [60] The underestimation of the PES is particularly significant for high-energy atomic configurations, which limits its application to molecular dynamics simulations and rare-event samplings. Recent large data efforts[31, 61] and open-sourced universal MLIPs[62, 63] that include more off-equilibrium atomic configurations show promise in addressing these challenges as discussed in Section 3.2. However, direct transferability of universal MLIPs for specific downstream molecular or materials modeling remains uncertain, as they are not yet demonstrably better than models built directly for the new system or property from scratch. We return to this issue in Section 3.5.
Another limitation is the lack of incorporation of long-range interactions. Although short-range potentials may be sufficient to describe most properties of homogeneous bulk systems [64], they may fail for interfaces [65], dielectric response [66, 67], and dilute ionic solutions with Debye-Hückel screening. For example, consider two charged molecules that are separated by a distance greater than the cutoff radius used in most MLIPs (Fig.1b). These models, which focus on short-range interactions, will fail to accurately describe the electrostatic forces between the molecules (Fig.1c). This issue cannot be resolved simply by using message passing, as the molecules exist on separate graphs and do not interact within the framework of short-range biases inherent to message-passing neural networks. On the other hand, MLIPs with built-in long-range corrections [68, 69, 70] can effectively solve such artifacts (Fig.1d).
Figure 1: (a) Distribution of softening scales ($f_{\text{MLIP}}/f_{\text{DFT}}$) sampled from 1,000 high-energy configurations using the structures from the WBM dataset.[57] Reproduced with permission from Ref. [60].(b) An example of two charged molecules that are far apart and do not interact due to (c) short-range implicit bias from GNNs and message passing, and (d) a new ML model that incorporates long-ranged information. [69]. Blue points are for training and red points are for testing. Adapted with permission from Ref. [69] (e) Depiction of instability in molecular simulations with MLIPs, which can occur even at very small time discretizations [71]. Adapted with permission from Ref. [71]
@@ 3.2 Data Requirements, Pitfalls, and Challenges
Regarding elemental diversity, most plane-wave DFT datasets follow the standards set by the Materials Project[33] and thus cover most of the periodic table up through plutonium. In contrast, molecular force field and DFT datasets are considerably more constrained in their elemental coverage, with most datasets only including C, H, N, and O and some coverage of S, F, Cl, and Br. The only molecular datasets of over 1 million data points with both energy and forces labels and which go beyond this set of elements are SPICE[78] (P, Li, Na, Mg, K, Ca, and I), AIMNet2 (P, I, Si, B, As, and Se), and more recently QCML (73 elements total)[79] and OMol25 (83 elements)[80]. The only large-scale datasets with variable-charge species are SPICE(2)[78], AIMNet2[81], AIMNet-NSE[82], solvated protein fragments[83]/GEMS[84], and most recently QCML[79] and OMol2025[80]. While AIMNet-NSE[82] is a dataset with variable spin species, it carries no force labels, whereas QCML does include both labels for some limited spin species[79] whereas variable spin is much better covered by OMol2025[80]. Another consideration for chemical diversity is the variety of chemical interactions captured by the data. While most of the molecular datasets mentioned thus far focus on isolated molecules, SPICE[78] makes an effort to capture varied intermolecular interactions between molecules, including some metal ions, solvated protein fragments[83]/GEMS[84] includes protein fragments with multiple nearby solvent species, and ANI-2x[85] includes dimer scans and water clusters. Recognizing the value of intermolecular interactions, OMol25 also includes these data types.[80]
In order to create structural diversity, large-scale data-acquisition efforts use a collection of sampling techniques such as normal modes, ab initio and empirical molecular dynamics, minimization and relaxation, reaction path sampling, metadynamics, and active learning, as outlined in Table 1. Structural diversity also encompasses the nature of the regions of the PES covered by the dataset. Having a broad range of forces between 0 eV/Å and $\sim$10 eV/Å is important to effectively learn chemically-critical components of the PES encountered during MD, geometry optimization, or reaction path and transition state optimization. To that end, Transition-1x is the only large-scale dataset which explicitly samples reaction paths and transition states – critical areas for any truly foundational MLIP.[86, 87] Extremely high
@@
OMol25[80] 100M $\omega$B97M-D3/def2-TZVPD Molecular dynamics
Rattling
Geometry optimization
Reaction pathways
QCML[79] 33.5M PBE0-D4/NAOs[88] Chemical graphs
Normal mode
@@
ANI-1x[92]/ANI-2x[85] 8.9M $\omega$B97X/6-31G* Dimer and torsion scans
@@
Alexandria[97] 30M PBE(+U)/PAW, SCAN/PAW Relaxation trajectory
@@
forces (>50 eV/Å) can be detrimental to training and are usually filtered out, although high-energy and force data can be weighted with a Boltzmann factor[16], which still provides the model important chemical information about these regions. Finally, given that LLM FMs are pre-trained in an unsupervised manner on unlabeled data, it is worth asking what sort of data would be required to realize an analogous procedure in the chemical context. Ji and co-workers recently demonstrated that self-supervised pre-training on 800 million unlabeled 3D conformations of organic molecules improved down-stream performance on QM9 property prediction.[75]
@@
Figure 2: Proposed areas where substantial additional high-quality DFT data should be generated in order to realize a general and transferable foundational MLIP. a) Beyond-organic elements. Reproduced with permission from Ref. [100]. b) Non-covalent interactions. Reproduced with permission from Ref. [101]. c) variable charge/spin/excited states. Adapted with permission from Ref.[102, 103]. d) Complex reactivity. Reproduced with permission from Ref.[104]. e) long-range interactions. Reproduced with permission from Ref. [105]. f) Solid-solid interfaces. Reproduced with permission from Ref. [106]. g) Surfaces under applied potential. Reproduced with permission from Ref. [107]. h) Defects. Reproduced with permission from Ref. [108]. i) Solid-liquid interfaces. Reproduced with permission from Ref. [109]. j) Rich quantum information. Left: Reproduced with permission from Ref. [110]. Right: Reproduced with permission from Ref. [111].
The specific areas where we believe the community should devote major effort to large-scale data generation are shown in Figure 2. For molecular DFT, the majority of the periodic table remains poorly covered, and under-representation of metal-organic complexes, complex electrolytes, metalloenzymes, and disordered proteins underscores the need for substantial dataset generation efforts beyond organic elements (Fig. 2a). These types of systems could also address three other major gaps: non-covalent interactions (Fig. 2b), systems where long-range interactions are critical (Fig. 2g), and variable charge, spin, and excited states[102, 103] (Fig. 2c) which have been poorly covered by current datasets. Excited state energies and properties also offer an avenue for dataset development, where time-dependent DFT (TDDFT) calculations[112] are relatively inexpensive (roughly the cost of ground state DFT per state) and have been extensively benchmarked[102] against higher level theories[113](Fig. 2e). We suggest that, beyond open-shell metals, combustion chemistry and atmospheric chemistry are spaces where different spin states are important and could be valuable. The final major gap we will describe is reactivity, particularly involving complex mechanisms and reactions in the condensed phase (Fig. 2d).
In the context of plane-wave DFT, we believe that major gaps include solid-solid interfaces (Fig. 2h), solid-liquid interfaces (Fig. 2j), surfaces under different applied potentials, (Fig. 2g), and defects (Fig. 2i). We note that some of these areas will require the use of better density functionals beyond the ubiquitous PBE and RPBE. Analogous to spin states for molecular properties, additional data generation for magnetic properties could also aid in correct magnetic moment initialization for diverse materials.
@@
Finally, we believe there is a major need for large-scale datasets of rich quantum information, such as electron densities, orbital information, QTAIM descriptors, NBO interactions, and beyond (Fig. 2k). We note that infusing rich quantum information into ML models has recently been shown to be beneficial[114, 110, 115, 111], and believe that powerful FM capabilities could be unlocked with such data. For example, MLIP FMs could be pre-trained on this extra information by swapping prediction schemes to predict atom and bond-level values. This could pave the way for foundational chemical models that go beyond MLIPs, serving as powerful representation models that can be adapted to predict forces and energies downstream.
@@
Going beyond DFT is desirable. Here it is appropriate to highlight the effort of Smith and co-workers in generating the ANI-1ccx data set containing energies and forces of 500,000 organic molecules obtained with an accurate CCSD(T)*/CBS composite extrapolation.[120, 92]. This protocol has deviations from the Schrödinger limit that are larger than normally acceptable for small, very high accuracy datasets,[126] but are much smaller than the best DFT errors, even in large basis sets.[2] Continued diverse collections of gold-standard CCSD(T) data (or beyond when necessary) would be extremely valuable for fine-tuning. Another result worth highlighting is the CCSD(T) assessment[127] of over 10,000 barrier heights in RDB7, which led to unexpectedly large RMS errors from the normally reliable $\omega$B97X-D3 functional. The largest of these deviations were resolved by removing orbital instabilities and allowing the DFT to be spin-polarized when necessary.[128] However this highlights the limitations of even hybrid DFT for reactive systems where strong correlation effects are at play, as revealed by such symmetry-breaking.[129] For the most part, CC benchmarks for condensed matter periodic systems lie in the future, although there has been notable progress in CC methods for materials science.[130]
The third factor are other error sources that prevent a given “model/basis” combination from yielding a well-defined value. In particular, DFT calculations on materials require careful attention to best practices[131, 125] to ensure reproducibility, beginning with choice of cutoffs, quadrature quality, and convergence criteria. Such factors have been discussed and investigated for the OC20 data.[132] There are also non-trivial additional sources of error. Reaching the thermodynamic limit is particularly challenging for hybrid functionals.[133] Pseudopotential errors, typically not present in molecular calculations involving light elements, also require careful evaluation[134, 135, 136]. Thresholds and cutoffs also affect molecular calculations, but can be controlled, and combined with SCF algorithms that ensure descent.[137] The local correlation methods needed for CCSD(T) on large molecules[138] are also subject to domain errors.[139] They scale with size, and may impact accuracy[140, 141] and so ideally should be removed and replaced by strict numerical thresholds.[142, 143]. These systematic errors make dataset fusion difficult as machine learning models trained on such datasets will likely have difficulty learning physical properties/labels from disjoint underlying distributions. In addition, reviewers could help find inconsistencies before initial publication and scientists remain committed to maintaining datasets after publication. One recent example underscoring these realities was a study by Garrison et. al.[144] where the authors trained machine learning models on the tmQM[145] dataset but found many structures in the dataset to contain missing hydrogen atoms. After filtering and recomputing the dataset, model performance improved significantly.